OpenStudy (anonymous):

Derivative y=x^cos(x) I think i am doing something completely wrong or just my steps... wondering if anyone can show step?

OpenStudy (tkhunny):

Let's see your steps. \[ln(y) = \cos(x)\cdot\ln(x)\] Now what?

OpenStudy (anonymous):

Implicit differentiation?

OpenStudy (anonymous):

you can do what @tkhunny suggests, but it is also important to know that the definition of \(b^x\) is \(e^{x\ln(b)}\) so you are looking at \[e^{\cos(x)\ln(x)}\] and you take the derivative using the chair rule you get \[e^{\cos(x)\ln(x)}\times \frac{d}{dx}[x\cos(x)]\]

OpenStudy (tkhunny):

Implicit and Logarithmic!

OpenStudy (anonymous):

if you say "take the log" once you find the derivative of \(\cos(x)\ln(x)\) which is where all the work is , you multiply by the original function also i made a typo, it should be \[e^{\cos(x)\ln(x)}\times \frac{d}{dx}[\cos(x)\ln(x)]\]

OpenStudy (anonymous):

i think i got a mistake on your step hunny, which is what i did.. but it's probably just me

OpenStudy (tkhunny):

Note: You DO have to be careful with such functions when they meet calculators and computer math programs (Maple, Mathematica, MatthCad to name a few). You may get only positive values for 'x' unless you try hard to get the negative ones. In full agreement with @satellite73, we shoudl also keep in mind that the transformation \[\log_{b}(a) = c \iff b^{c} = a\] for suitable a, b, and c is quite generally applicable. This makes my logarithmic form and satellites73 exponential form very much the same. Just wade carefully thougth the chain rule, the product rule, and a little algebra.\[\ln(y) = cos(x)\cdot\ln(x)\]leads to \[\frac{y'}{y} = \frac{\cos(x)}{x} -\sin(x)\cdot\ln(x)\]